Easter Egg and Poultry Company has $1,190,000 in assets and $647,000 of debt. It reports net income of $185,000. Having a hard time finding part b of the question below. What is its return on stockholders’ equity?
Answer #1

Total assets=debt+equity

equity=(1,190,000-647,000)=$543,000

ROE=net income/equity

=185,000/543,000

=34.07%(Approx).